i need help with the problem: -2/3(6-2a)=6
Answer by MathPro(14)

Good Afternoon,
First expand the equation, by multiplying through by (2/3):
-(2/3)6 + (2/3)2a=6
Simplify:
-12/3 + 4a/3 = 6
-4+4a/3=6
Solve for a:
(1)move the 4 to the right-hand-side by adding 4 to each side:
-4 +4 +4a/3=6+4
4a/3=10
(2) move the fraction to the right-hand-side too by division
(3/4)(4a/3)=10(3/4)
a=30/4
simplify the fraction:
a=15/2
check the answer:
substitute a=15/2 into the original equation:
-(2/3)6+(2/3)2(15/2)=6
-4+10=6
The left-hand-side is equal to the right-hand-side so the solution is correct
Good Luck!
Selma from Rent-A-Whiz, an online tutoring and homework help website. Just visit www.rentawhiz.com to learn more. Currently, we are offering to solve one math problem for free. Just visit our website, select the homework tab, then click on the math problem submission link and quote coupon code 003 when you submit your problem.